OpenStudy (anonymous):

What is the slope of the line that passes through (-4, 5) and (2, -3)?

OpenStudy (anonymous):

the question different from question before

OpenStudy (anonymous):

take (x1,y1) =(-4,5) (x2,y2) =(2,-3) m = (y1-y2)/(x1-x2) m = -(4/3)

OpenStudy (anonymous):

tanks